OpenStudy (anonymous):

find the double integral the integral from 5 to 4 the integral from 5 to 4 (4x+y)^-1

OpenStudy (reemii):

- integrating \(\frac1{4x+y}\) with respect to \(y\) gives \(\ln(4x+y)\). - a primitive of \(\ln(x)\) is \(x\ln(x)- x \). This might help, I didn't do it but im confident it's what you need.

OpenStudy (anonymous):

im not really sure what you did im just had my first lecture on double integrals today so im not sure how to solve it

OpenStudy (anonymous):

Marry me, I'm an Asian.

OpenStudy (reemii):

you can choose the order of integration. I'll fix it to this: \(\int_4^5(\int_4^5 \dots\,dy)dx\) You will first evaluate what's inside the parenthesis.

OpenStudy (reemii):

since it's an integration of the variable \(y\) you consider \(x\) as constant. A primitive of the function is ln(4x+y). -> the parenthesis is equal to \(\ln(4x+5) - \ln(4x+4)\). Then you can split the integrals and use "normal" integration.

OpenStudy (anonymous):

so the equation you typed is my function and i first have to find the integral with respect to y and whatever i get from there i take the integral with respect to x

OpenStudy (reemii):

yes. in general, the parenthesis , after integrating once, still contains one variable. Here \(dy\) is in the parenthesis. After comupting this, all \(y\)'s have disappeared. only \(x\)'s remain.

OpenStudy (anonymous):

ok so the integral for the above functions is 1/4x+5 - 1/4x+4 right

OpenStudy (reemii):

which integral? I don't think this function appears in the answer. One part of the answer is \[\int_4^5 \ln(4x+5) \,dx = \int_{21}^{25} \ln(u) \frac14\,du\\ = \left[ u\ln u - u \right]_{21}^{25} \]i think.

OpenStudy (dan815):

hi im here!!

OpenStudy (reemii):

i forgot \(\frac14\)

OpenStudy (dan815):

my favourite questions

OpenStudy (dan815):